OpenStudy (cutiekawdfi):

help? determine the discriminant. how many real solutions does it have? 3x^2 + 3 = 6x

OpenStudy (anonymous):

what is this ^

pooja195 (pooja195):

We first need to put it in standered form \[\huge~\rm~ax^2+bx+c=0\]

OpenStudy (cutiekawdfi):

how do i do that

pooja195 (pooja195):

Start by subtracting 6x from both sides

pooja195 (pooja195):

\[\huge~\rm~3x^2 + 3 -6x= 6x -6x\]

OpenStudy (cutiekawdfi):

3x^2+ -3x?

pooja195 (pooja195):

Nope the 6x becomes part of the equation so you get \[\huge~\rm~3x^2 -6x+3= 0\]

OpenStudy (cutiekawdfi):

why does it though

pooja195 (pooja195):

What do you mean?

pooja195 (pooja195):

It always was we just made our equation equal 0 since its a quadratic expression we need it in standered for which is \[\huge~\rm~ax^2+bx+c=0\] from there we can name our abc values \[\huge~\rm~3x^2 -6x+3= 0\] a=3 b=-6 =c=3 Do you understand?

OpenStudy (cutiekawdfi):

okay

OpenStudy (cutiekawdfi):

@pooja195

pooja195 (pooja195):

y=3x2−6x+3 discriminant=\[\huge~\rm~b^2−4(a)(c)\]

pooja195 (pooja195):

discriminant=\[\huge~\rm~(-6)^2−4(3)(3)\] Solve